# 3D ISM structure challenges the Serkowski relation

### v1

@@ 4.3 On the parameters of the Serkowski formula
Here we demonstrate how the derived parameters of the Serkowski relation can be incorrect, when erroneously assuming a single cloud in the LOS. We explored how 3D effects affect the distributions of $\lambda_{max}$, and $P_{max}$ that one obtains through the classical fitting of the Serkowski relation in the $P-\lambda$ space; $K$ linearly correlates with $\lambda_{max}$ (Wilking et al., 1980), and thus can be omitted from this analysis. We fit the data only in $P-\lambda$ space (Eq. 1) by minimizing
@@
assuming a one-cloud model, and imposing the same priors as discussed in Sect. 3.3. Since the effect of the 3D dust structure is more prominent in $\theta-\lambda$ space, fits in $P-\lambda$ were comparably good for all four subsamples. We then explored the normalized distributions of the derived parameters for all subsamples (Fig. 5 for $\lambda_{max}$, Fig. 6 for $P_{max}$).
Past studies of the Serkowski parameters show that the distribution of $\lambda_{max}$ is singly-peaked, with an average value close to 0.5 $\mu$m (Martin & Whittet, 1990). In contrast, we find that for $N_{C}=1$ (single-cloud sightlines) the intrinsic distribution of $\lambda_{max}$ is bimodal, with peaks at approximately 0.28 and 0.62 $\mu$m. On the other hand, all distributions with $N_{C}>1$ (multiple-cloud sightlines) are unimodal and peak at around 0.55 $\mu$m, which is close to the value that is considered as the Galactic average. By performing Monte Carlo (MC) simulations we found that the unimodality in the $\lambda_{max}$ distribution, which is observed for $N_{C}>1$, can be reproduced with multiple-cloud models with values of $\lambda_{max}$ randomly drawn from the $N_{C}=1$ distribution. The intrinsic Galactic distribution of $\lambda_{max}$ may thus be bimodal, and the currently accepted value for the Galactic average of $\lambda_{max}$ ($\sim 0.5\mu$m) is likely contaminated by LOS effects. Notably, the larger spread of the distribution of $\lambda_{max}$ for the $N_{C}=2$ case is not surprising. Clarke & Al-Roubaie (1984) had theorized that in the case of two clouds in the LOS, the acquired $\lambda_{max}$ from the Serkowski fit is different from the average $\lambda_{max}$ of the two clouds when the two clouds have different parameters. Most areas containing stars behind a single cloud also contain stars behind multiple clouds. We calculated the average $\lambda_{max}$ for targets in distinct areas on the sky (Fig. 7). In most cases, the average $\lambda_{max}$ differs significantly between subsets (see Table 3) located in the same area. This suggests that using the $\lambda_{max}$ of a star from an area with $N_{C}>1$ as a characteristic value for the LOS dust properties could be misleading.
@@
Figure 6: Normalized distributions of $P_{max}$ for subsamples with different number of clouds in the LOS. See Sect. 4.3 for details.
Table 3: Mean $\lambda_{max}$ values in $\mu m$ and their error on the mean for different regions in the sky and different subsamples. Regions with N/A in the error of the mean are cases where there is only one star in the specific region for the subsample.
@@
Figure 7: Positions of targets depending on their $N_{C}$. Maps are in Galactic coordinates, in mollweide projection, centered in (0,0). Parallels are drawn every 30o. Meridians are drawn every 60o. Longitude increases towards the right-hand side. Colors correspond to the average $\lambda_{max}$ of each area. Marked areas with ’A’ through ’G’ contain stars that are located behind a single as well as multiple clouds. See discussion in Sect. 4.3.
The distribution of $P_{max}$ is also affected by the number of clouds along the LOS. We can immediately see this by assuming, without loss of generality, that two clouds lie along a LOS. The polarization induced by the two clouds can either (partially or totally) add up, or (partially or totally) cancel out, depending on the difference between the EVPAs each would impart on its own.When the EVPA difference is close to 0o the polarization adds up, while when the difference is close to 90o it cancels out. This is exactly what we observe in the distributions of $P_{max}$ (Fig. 6). The $P_{max}$ distribution for $N_{C}=1$ is free from LOS effects, hence it should resemble the ”ground truth”. This distribution is unimodal and peaks at $\sim 2.5\%$. On the other hand, the $N_{C}=2$ distribution is bimodal, with both peaks systematically offset from the mode of the $N_{C}=1$ distribution. The bimodality of the $N_{C}=2$ distribution can be understood by the LOS integration effects acting constructively (generating the high-$P_{max}$ peak), or destructively (generating the low-$P_{max}$ peak). The distribution for $N_{C}=3$ displays similar peaks, though it is smoother, as the addition of a third cloud facilitates producing polarization closer to the average observed in the $N_{C}=1$ case. Similarly to $N_{C}=1$, the distribution of the $N_{C}>3$ case is unimodal because an increased number of integration components naturally leads to a distribution with a well-defined peak, as a consequence of the central limit theorem. Overall, both the $\lambda_{max}$, and the $P_{max}$ are significantly affected by the 3D structure of the ISM, which, if ignored, can lead to erroneous Serkowski best-fit parameters.
@@
Our results suggest that fitting a single Serkowski relation is invalid in LOSs with multiple clouds. There can be cases where a single Serkowski relation yields good fits, even when 3D effects are important (Sect. 4.1). However, in these cases the best-fit parameters of a single Serkowski curve are misleading, and not representative of the dust grain physics.
@@ 5.1 Implications for dust modeling
The findings presented in this study have significant implications for dust modeling. Dust models, and grain alignment theories must satisfy the constraints obtained from Serkowski fits (Martin & Whittet, 1990; Andersson et al., 2015; Draine, 2024). However, our results indicate that the Serkowski parameters derived from fitting the observed polarization data in complex interstellar environments may not accurately represent the underlying dust properties, leading to erroneous conclusions about grain size distributions and composition. The results from Sect. 4.3 suggest that the intrinsic distribution of $\lambda_{max}$ in the clouds of our Galaxy may actually be bimodal (Fig. 5). This finding contradicts past studies (Martin & Whittet, 1990), which likely stem from (incorrectly) fitting the Serkowski formula in regions with multiple clouds. Although, our employed sample is the largest to-date, the statistics are still limited, and more data are required to further explore the intrinsic distribution of $\lambda_{max}$ in our Galaxy.
@@ 5.2 Implications for the identification of intrinsically polarized stars
It is customary to iterpret deviations from the Serkowski relation in the $P(\lambda)$ behavior of point sources as indicative of intrinsic polarization mechanisms, such as the presence of circumstellar disks (e.g. Topasna et al., 2023). However, the wavelength dependence of polarization in such cases is known to lack a typical pattern (Bastien, 2015). In contrast, we have demonstrated that fitting simple two-cloud models can account for deviations from the Serkowski formula quite effectively (Sect. 4.1). This suggests that deviations from the Serkowski curve alone are insufficient to indicate the presence of an intrinsic polarization component. In order to validate that a star not following the Serkowski formula is indeed intrinsically polarized, polarization data could be combined with other evidence, such as detection of significant variability, or comparison with neighboring stars.
@@ 5.3 The possibility of 3D tomography with the Serkowski relation
Fitting multiple-cloud models to multiwavelength (optical to near-infrared) polarization data paves the way to constraining the LOS variations of the plane-of-the-sky magnetic field morphology (tomography). If the number of clouds in the LOS is known, for instance from 3D extinction maps (e.g., Edenhofer et al., 2024; Green et al., 2019), it is possible to extract the parameters of each individual cloud using the techniques discussed in Sect. 3.4. However, the following considerations should be kept in mind. Firstly, it is possible that two or more clouds in the LOS have similar parameters. In such cases, the polarization data may be adequately fit with fewer clouds than what the 3D extinction maps suggest. For example, if two clouds along a LOS have comparable $\lambda_{max}$ and $\theta$, their polarization contributions will combine, resulting in a profile resembling that of a single cloud but with an increased degree of polarization. This might explain our target star HD93222 (Sect. 4.1), where the extinction profile displays at least three distinct steps, yet the data are sufficiently well-fit by a 2-Cloud model. Another possibility is that the third cloud has a low abundance of aligned grains.
Another strategy one could follow to identify cases of clouds with comparable properties is to obtain measurements of multiple stars at different distances along the same LOS. The polarization signal would vary with distance, as each star would only be influenced by the foreground dust column. Thus, stars lying behind the same clouds would show similar polarization trends, while distinct changes would only occur for stars behind different clouds. Following such methods could complement the BISP-1 algorithm (Pelgrims et al., 2023), which employs single-band optical polarization data with parallaxes from Gaia, for performing ISM tomography along many LOSs (Pelgrims et al., 2024) in the context of optopolarimetric surveys, such as the Pasiphae survey (Tassis et al., 2018). Thus, multiwavelength polarization data promises to provide an independent method for performing ISM magnetic field tomography, which could usher in the era of high-precision 3D magnetized ISM cartography.
@@ 6 Summary and Conclusions
In this paper, we showcase the effect of the 3D structure of the ISM on the well-known and widely-used Serkowski relation. Our findings are summarized as follows.
@@
    We revisited theoretically how the Serkowski formula could become invalid in LOSs with many dust components. We showed that the variations of both the degree of polarization and EVPA with wavelength can be severely affected in some stars. In other stars, the polarization profile may misleadingly appear to follow the Serkowski relation, despite the contribution from multiple dust clouds. In such cases, the imprint of the 3D structure can be mainly observed in the $\theta-\lambda$ profile.
@@
    As a test case, we selected two different targets that appear to be behind multiple dust clouds, based on their extinction profiles, and for which archival spectropolarimetric data exist. We demonstrated that the single-cloud model, implicitly assumed for Serkowski fits, fails in these cases, while we successfully fitted 2-cloud models.
@@
    We proposed the $q-u$ fitting technique as the most appropriate approach to fit the Serkowski relation. This method allows for tracing both $P$ and $\theta$ as functions of wavelength simultaneously, enabling fitting for more than one clouds along the LOS.
@@
    We used a sample of 223 stars behind different number of clouds with polarimetric measurements in three or more bands, to fit and evaluate statistically the performance of the Serkowski formula. We found that the Serkowski formula is not a good model for LOSs with two or more clouds, while it fits well the data in the cases of one or more than three clouds in the LOS. The latter is an outcome of averaging.
@@
    Our analysis of the fitted parameters of the 223 stars suggests that the intrinsic distribution of the $\lambda_{max}$ parameter in our Galaxy may be bimodal, something that was veiled until now, due to treating sightlines with many clouds as a single dust component.
@@
    A poor Serkowski fit by itself does not constitute conclusive evidence of intrinsic polarization in a source.
@@
    Fitting the Serkowski formula for stars behind multiple dust clouds leads to an incorrect estimation of the fitted parameters, and, thus, of the underlying dust physics.
@@
    Our results hint towards the possibility of performing magnetic tomography using multiband polarization data in combination with Gaia distances.
In conclusion, the Serkowski relation should be cautiously applied when the ISM 3D structure is complex.

### v3

@@ 4.3 The parameters of the Serkowski formula
Here, we demonstrate how the derived parameters of the Serkowski relation can be incorrect, when erroneously assuming a single cloud in the LOS. We explored how 3D effects affect the distributions of $\lambda_{max}$, and $P_{max}$ that one obtains through the classical fitting of the Serkowski relation in the $P-\lambda$ space; $K$ linearly correlates with $\lambda_{max}$ (Wilking et al., 1980), and thus can be omitted from this analysis. We fit the data only in $P-\lambda$ space (Eq. 1) by minimizing
@@
assuming a one-cloud model, and imposing the same priors as discussed in Sect. 3.3. For these fits, we used the debiased $P$ values. Since the effect of the 3D dust structure is more prominent in $\theta-\lambda$ space, fits in $P-\lambda$ were comparably good for all four subsamples. We then explored the normalized distributions of the derived parameters for all subsamples (Fig. 8 for $\lambda_{max}$, Fig. 9 for $P_{max}/A_{V}$).
Past studies of the Serkowski parameters show that the distribution of $\lambda_{max}$ is singly peaked, with an average value close to 0.5 $\mu$m (Martin & Whittet, 1990). In contrast, we find that for $N_{C}=1$ (single-cloud sightlines) the intrinsic distribution of $\lambda_{max}$ is bimodal, with peaks at approximately 0.28 and 0.62 $\mu$m. On the other hand, all distributions with $N_{C}>1$ (multiple-cloud sightlines) are unimodal and peak at around 0.55 $\mu$m, which is close to the value that is considered to be the Galactic average. By performing MCMC simulations, we found that the unimodality in the $\lambda_{max}$ distribution, which is observed for $N_{C}>1$, can be reproduced with multiple-cloud models with values of $\lambda_{max}$ randomly drawn from the $N_{C}=1$ distribution. The intrinsic Galactic distribution of $\lambda_{max}$ may thus be bimodal, and the currently accepted value for the Galactic average of $\lambda_{max}$ ($\sim 0.5\mu$m) is likely contaminated by LOS effects. Notably, the larger spread of the distribution of $\lambda_{max}$ for the $N_{C}=2$ case is not surprising. Clarke & Al-Roubaie (1984) theorized that in the case of two clouds in the LOS, the acquired $\lambda_{max}$ from the Serkowski fit is different from the average $\lambda_{max}$ of the two clouds when the two clouds have different parameters. Furthermore, it is worth considering that most areas containing stars behind a single cloud also contain stars behind multiple clouds. We calculated the average $\lambda_{max}$ for targets in distinct areas on the sky (Fig. 3). In most cases, the average $\lambda_{max}$ differs significantly between subsets (see Table 3) located in the same area. This suggests that using the $\lambda_{max}$ of a star from an area with $N_{C}>1$ as a characteristic value for the LOS dust properties could be misleading.
@@
Figure 9: Normalized distributions of $P_{max}$ for subsamples with different number of clouds in the LOS. See Sect. 4.3 for details. The following bins are omitted for clarity. For $N_{C}=2$: bin at $P_{max}/A_{V}\simeq 23$ with a height of 0.0135. For $N_{C}=3$: bin at $P_{max}/A_{V}\simeq 13$ with a height of 0.0145.
Table 3: Mean $\lambda_{max}$ values in $\mu m$ and their error on the mean for different regions in the sky and different subsamples.
@@
 [^3] Regions with N/A in the error of the mean are cases in which there is only one star in the specific region for the subsample.
The distribution of $P_{max}/A_{V}$ is also affected by the number of clouds along the LOS. We can immediately see this by assuming, without loss of generality, that two clouds lie along a LOS. The polarization induced by the two clouds can either (partially or totally) add up, or (partially or totally) cancel out, depending on the difference between the EVPAs each would impart on its own. When the EVPA difference is close to 0o, the polarization adds up, while when the difference is close to 90o it cancels out. On the other hand, the extinction, $A_{V}$, induced by each cloud is always additive. Therefore, with more clouds, we would expect to have in general lower $P_{max}/A_{V}$ than in the $N_{C}=1$ cases. However, since $P_{max}$ is a derived parameter rather than a direct measurement, the fit can be poor in cases with $N_{C}>1$, and $P_{max}$ may not represent any physical truth. In such scenarios, $P_{max}$, and thus $P_{max}/A_{V}$, could take on arbitrary values. This is exactly what we observe in the distributions of $P_{max}/A_{V}$ (Fig. 9). The $P_{max}/A_{V}$ distribution for $N_{C}=1$ is free from LOS effects, and hence it should resemble the “ground truth.” This distribution peaks at $P_{max}/A_{V}\simeq 3-3.5\%/$ mag, with only a few points with $P_{max}/A_{V}>4.5\%/$ mag. On the other hand, the $N_{C}=2$ distribution peaks at a lower value of $P_{max}/A_{V}\simeq 2\%/$ mag, with its tail extending to higher values of $P_{max}/A_{V}$ than when $N_{C}=1$. The shape of the $N_{C}=2$ distribution exactly follows the expectations discussed above. Similarly, for the distributions of more components, $P_{max}/A_{V}$ consistently has lower values than the case of $N_{C}=1$.
The polarization-to-extinction ratio has garnered significant interest within the ISM community (e.g. Hiltner, 1956; Serkowski et al., 1975; Panopoulou et al., 2019; Planck Collaboration et al., 2020; Angarita et al., 2023). While here we present our findings for $P_{max}/A_{V}$, discussions typically focus on $P/A_{V}$ in the context of measured polarization in a specific band, rather than on a fit parameter such as $P_{max}$. Consequently, our results for the distributions with $N_{C}>1$ cannot be fully integrated into this broader discussion if the fit parameter, $P_{max}$, does not accurately reflect the true value for a given field. On the contrary, the findings for the $N_{C}=1$ case should be free from bias. Within this subsample, we identify three targets with extreme $P_{max}/A_{V}$ values: two with $P_{max}/A_{V}\simeq 10-11\%/\mathrm{mag}$ and one with $P_{max}/A_{V}\simeq 6\%/\mathrm{mag}$. These values correspond to $P_{max}/E(B-V)\simeq 31-34\%/\mathrm{mag}$ and $P_{max}/E(B-V)\simeq 18.6\%/\mathrm{mag}$, respectively, assuming that $R_{V}=3.1$ with $R_{V}=A_{V}/E(B-V)$ (Fitzpatrick, 2004). These values exceed the highest $P_{max}/E(B-V)$ reported by Angarita et al. (2023) to date. However, the observed maximum $P/E(B-V)$ has consistently been revised upwards over time: from $P/E(B-V)\leq 9\%/\mathrm{mag}$ (Hiltner, 1956; Serkowski et al., 1975), to $P/E(B-V)\leq 13\%/\mathrm{mag}$ (Panopoulou et al., 2019; Planck Collaboration et al., 2020), and most recently to $P/E(B-V)\leq 16\%/\mathrm{mag}$ (Angarita et al., 2023). While our values may appear extreme, they may well reflect true physical conditions, and further research is warranted.
@@
Overall, both the $\lambda_{max}$ and the $P_{max}/A_{V}$ are significantly affected by the 3D structure of the ISM, which, if ignored, can lead to erroneous Serkowski best-fit parameters. Our results suggest that fitting a single Serkowski relation is invalid in LOSs with multiple clouds. There can be cases in which a single Serkowski relation yields good fits, even when 3D effects are important (Sect. 4.2). However, in these cases the best-fit parameters of a single Serkowski curve are misleading, and not representative of the dust grain physics.
All the targets used in the statistical sample are presented in Table 4 with their properties. The complete table can be found only in electronic form (Sect.Data availability).
Table 4: Targets and their properties.
Gaia ID RA Dec Distance $A_{V}$ $N_{bands}$ $N_{C}$ $\chi^{2}_{\nu}$ $P_{max}$ $K$ $\lambda_{max}$ $\theta$
[…]
180979373413769344 80.68056 33.28809 1151 0.61 4 ¿3 0.61 1.30 0.56 0.350 -16.7
 [^4] The full table is available at the CDS.
@@ 5.1 Implications for dust modeling
The findings presented in this study have significant implications for dust modeling. Dust models and grain alignment theories must satisfy the constraints obtained from Serkowski fits (Martin & Whittet, 1990; Andersson et al., 2015; Draine, 2024). However, our results indicate that the Serkowski parameters derived from fitting the observed polarization data in complex interstellar environments may not accurately represent the underlying dust properties, leading to erroneous conclusions about grain size distributions and composition. The results from Sect. 4.3 suggest that the intrinsic distribution of $\lambda_{max}$ in the clouds of our Galaxy may actually be bimodal (Fig. 8). This finding contradicts past studies (Martin & Whittet, 1990), which likely stem from (incorrectly) fitting the Serkowski formula in regions with multiple clouds. Although, our employed sample is the largest to date, the statistics are still limited, and more data are required to further explore the intrinsic distribution of $\lambda_{max}$ in our Galaxy.
@@ 5.2 Implications for the identification of intrinsically polarized stars
It is customary to interpret deviations from the Serkowski relation in the $P(\lambda)$ behavior of point sources as indicative of intrinsic polarization mechanisms, such as the presence of circumstellar disks (e.g. Topasna et al., 2023). However, the wavelength dependence of polarization in such cases is known to lack a typical pattern (Bastien, 2015). In contrast, we have demonstrated that fitting simple two-cloud models can account for deviations from the Serkowski formula quite effectively (Sect. 4.2). This suggests that deviations from the Serkowski curve alone are insufficient to indicate the presence of an intrinsic polarization component. In order to confirm that a star not adhering to the Serkowski formula is intrinsically polarized, other evidence, such as the detection of variability in the polarization or a difference in polarization between the star and its neighbors, is required.
@@ 5.3 The possibility of 3D tomography with the Serkowski relation
Fitting multiple-cloud models to multiwavelength (optical to near-infrared) polarization data paves the way for constraining the LOS variations in the plane-of-the-sky magnetic field morphology (tomography). If the number of clouds in the LOS is known, for instance from 3D extinction maps (e.g., Edenhofer et al., 2024; Green et al., 2019), it is possible to extract the parameters of individual clouds using the techniques discussed in Sect. 3.4. However, the following considerations should be kept in mind. It is possible that two or more clouds in the LOS have similar parameters. In such cases, the polarization data may be adequately fit with fewer clouds than what the 3D extinction maps suggest. For example, if two clouds along a LOS have comparable $\lambda_{max}$ and $\theta$, their polarization contributions will combine, resulting in a profile resembling that of a single cloud but with an increased degree of polarization. This might explain our target star HD93222 (Sect. 4.2), where the extinction profile displays at least three distinct steps, yet the data are sufficiently well fit by a 2-Cloud model. Another possibility is that the third cloud has a low abundance of aligned grains, because of, for example, a magnetic field with a direction along the LOS.
Another strategy one could follow to identify cases of clouds with comparable properties is to obtain multiband or, ideally, spectropolarimetric measurements of multiple stars at different distances along the same LOS. The polarization signal would vary with distance, as each star would only be influenced by the foreground dust column. Thus, stars lying behind the same clouds would show similar polarization trends, while distinct changes would only occur for stars behind different clouds. Following such methods could complement the BISP-1 algorithm (Pelgrims et al., 2023), which employs single-band optical polarization data with parallaxes from Gaia. This algorithm is particularly relevant for performing ISM tomography along many LOSs (Pelgrims et al., 2024), especially in the context of optopolarimetric surveys, such as the Pasiphae survey (Tassis et al., 2018). While the method described by Pelgrims et al. (2023) can determine the number of clouds along a LOS and their distances, it does not fully capture the dust physics within each cloud. Multiband polarimetric measurements of a few stars at varying distances along the LOS can provide insights into the dust physics within each cloud. Thus, multiwavelength polarization data promises to provide an independent method for performing ISM magnetic field tomography, which could usher in the era of high-precision 3D magnetized ISM cartography.
@@ 6 Summary and conclusions
In this paper, we showcase the effect of the 3D structure of the ISM on the well-known and widely used Serkowski relation. Our findings are summarized as follows.
@@
    We have revisited theoretically how the Serkowski formula could become invalid in LOSs with many dust components. We have shown that the variations in both the degree of polarization and EVPA with wavelength can be severely affected in some stars. In other stars, the polarization profile may misleadingly appear to follow the Serkowski relation, despite the contribution from multiple dust clouds. In such cases, the imprint of the 3D structure can mainly be observed in the $\theta-\lambda$ profile.
@@
    As a test case, we selected two different targets that appear to be behind multiple dust clouds, based on their extinction profiles, and for which archival spectropolarimetric data exist. We demonstrated that the single-cloud model, implicitly assumed for Serkowski fits, fails in these cases, while we successfully fit two-cloud models.
@@
    We propose the $q-u$ fitting technique as the most appropriate approach to fit the Serkowski relation. This method allows one to trace both $P$ and $\theta$ as functions of wavelength simultaneously, enabling fitting for more than one cloud along the LOS.
@@
    We used a sample of 223 stars behind different numbers of clouds with polarimetric measurements in three or more bands, to fit and evaluate statistically the performance of the Serkowski formula. We found that the Serkowski formula is not a good model for LOSs with two or more clouds, while it fits the data well in the cases of one or more than three clouds in the LOS. The latter is an outcome of averaging.
@@
    Our analysis of the fit parameters of the 223 stars suggests that the intrinsic distribution of the $\lambda_{max}$ parameter in our Galaxy may be bimodal, something that was veiled until now, due to treating sightlines with many clouds as a single dust component.
@@
    A poor Serkowski fit does not by itself constitute conclusive evidence of intrinsic polarization in a source.
@@
    Fitting the Serkowski formula for stars behind multiple dust clouds leads to an incorrect estimation of the fit parameters, and thus of the underlying dust physics.
@@
    Our results hint at the possibility of performing magnetic tomography using multiband polarization data in combination with Gaia distances.
In conclusion, the Serkowski relation should be cautiously applied when the ISM 3D structure is complex. Finally, our findings underscore the critical role of spectropolarimetry in driving future advancements in ISM research. While wide-field polarimetry offers valuable insights, its limitations become evident when considering the physics of dust. Therefore, we strongly advocate for the widespread adoption of spectropolarimetry within the ISM community for future research.
Data availability
Table 4 is only available in electronic form at the CDS via anonymous ftp to cdsarc.u-strasbg.fr(130.79.128.5) or via http://cdsweb.u-strasbg.fr/cgi-bin/qcat?J/A+A/